# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

### v2

@@ V Neuronal correlations shape the sublinear scaling of memory capacity
It is noteworthy that the $n$-th term corresponds exactly to the $n$-th term in the Neumann series expansion, Eq.(5). Therefore, under the assumption of vanishing neuronal correlations ($\tilde{C}=O$), the summation terms in Eq.(V) vanish, resulting in $r(L)=1$. This implies that higher neuronal correlations contribute to the faster decay of $r(L)$ through the contributions of the summation terms in Eq.(V). In fact, we have confirmed that $r(L)$ decays more rapidly for hyperparameters that lead to higher neuronal correlation, such as larger input signals ($\tilde{\sigma}_{s}^{2}$), smaller recurrent weights ($g$), and lower noise intensity ($\sigma_{n}^{2}$), as depicted in Fig.3. Notably, while the slope of the memory capacity curve at the origin, which is proportional to $MC(1)$, reaches its maximum when the reservoir RNN operates near the edge of chaos [8, 9], the extent of the decay of $r(L)$ varies monotonically with the hyperparameters $g$, $\sigma_{s}^{2}$ and $\sigma_{n}^{2}$, irrespective of whether the RNN is near the edge of chaos. These results suggest that instead of proximity to the edge of chaos, neuronal correlation is a novel mechanism determining the computational ability of RC.
@@
Thus far, we have derived the memory capacity under the specific scaling assumptions (Eq.(3)). However, practical applications often deviate from these conditions, raising the question of whether neuronal correlations remain a key determinant of sublinear memory capacity growth in more general settings. To address this, Fig.4 explores a broader range of model parameters including larger input intensities and various noise levels, and different values of $g$. Here, we introduced a half-life of memory capacity growth, $L_{\rm half}$, defined as the value of $L$ at which $r(L)$ equals $0.5$ (Fig.4, inset). A lower $L_{\rm half}$ indicates more rapid decay of $r(L)$. Notably, as shown in Fig.4, strong neuronal correlations are closely associated with a faster decay of memory capacity growth, highlighting their persistent influence even when the original scaling assumptions are relaxed.
Figure 4: Relation between the level of neuronal correlations and the half-life of memory capacity growth. Each data point corresponds to a set of hyperparameters sampled from $g\sim U(0.2,2)$, $\sigma_{s}\sim U(0.1,3)$ and $\sigma_{n}\sim U(0,3)$ (100 samples are plotted). Red points indicate parameter regimes beyond our theoretical framework, as determined by $L_{\rm half}$ surpassing the deviation threshold of the theoretical values of $MC$ (Fig.2). For each parameter set, the half-life of memory capacity growth, $L_{\rm{half}}$, and pair-wise averaged absolute values of Pearson correlation coefficients of neuronal activity, $\rho_{ij}$, are calculated across 10 network and input realizations. The mean values obtained from these 10 realizations are displayed as a single plot in the figure. The activation function is $\phi(x)=\tanh(x)$. The network size is $N=1000$, and the simulation time length is $T=10^{4}$.
VI Scaling behavior of nonlinear computational capability
@@
While our previous analysis has focused on memory capacity, non-linear computational ability is equally crucial for RC to address complex tasks. As established in previous works, there exists a trade-off between memory capacity and nonlinear computational ability for RC  [7, 21]. Thus, we expect that nonlinear computational ability increases instead of a deceleration in the growth rate of memory capacity.
@@
Information processing capacity (IPC) introduced by Dambre et al. [7] offers a task-independent metric for evaluating the performance of RC, enabling a comprehensive assessment of both linear and non-linear computational capabilities. Since IPC theory is somewhat complicated, we provide detailed explanations in Appendix.C.1. Put simply, the IPC for degree $D$, denoted by $IPC_{D}$, represents the reservoir’s ability to approximate $D$-th order polynomial functions of past white noise input signals. The IPC has two key properties. First, by definition, memory capacity is exactly equivalent to $IPC_{1}$, and thus, $IPC_{D}$ for $D\geq 2$ represents non-linear computational ability. Second, it has been proven [7] that the total IPC, defined as $IPC_{\rm total}\equiv\sum_{D=1}^{\infty}IPC_{D}$, equals to the number of readout units, $L$, provided that the reservoir satisfies the echo state property, a necessary condition for a reservoir to be usable for RC. Here, the echo state property implies that the reservoir’s state is uniquely determined solely by input signals regardless of reservoir’s initial state [2].
@@
Fig.5 illustrates how the values of $IPC_{D}$ for our RNN model (Eq.(1)) change with increasing $L$ across various hyperparameter combinations ($g$, $\sigma_{s}^{2}$, and $\sigma_{n}^{2}$), as obtained through numerical simulations. Note that, due to the symmetry of the model, $IPC_{D}$ values for even-degree $D$ are identically zero. As shown, a decline in the growth of the $IPC_{1}$ (equivalent to memory capacity) is accompanied by an increase in the non-linear computational ability represented by $IPC_{D}$ for $D\geq 2$, reflecting memory-nonlinearity trade-off.
Our novel finding is that, as $L$ increases, successively higher-degree IPCs are sequentially activated from lower to higher orders, accompanied by supralinear rises in the reservoir’s nonlinear computational capabilities. This phenomenon is observed not only in an RNN satisfying the echo state property (Fig.5(a)), but also in a chaotic RNN (Fig.5(b)) and in an RNN subject to neuronal noise (Fig.5(c)), where the echo state property breaks down. Additionally, the same finding is observed for an RNN employing the ReLU activation function (Fig.8), where our theoretical framework for deriving memory capacity is not applicable (see Appendix.C.2).
@@
As analytical derivation of higher-degree IPC is beyond our capability, the detailed mechanism underlying these observations remains unclear. However, it is plausible that the neuronal correlation plays a pivotal role, analogous to its influence on memory capacity. This hypothesis is supported by speculation that each $IPC_{D}$ would exhibit linear scaling with $L$ if the effects of neuronal correlations were neglected, which contrasts with the observed supralinear and sequential emergence of $IPC_{D}$ (see Appendix.C.3).
Figure 5: IPC values for varying the number of readout neurons $L$ of a reservoir RNN with activation function $\phi(x)=\tanh(x)$ and network size $N=1000$. The top row panels show stacked representations of IPC values for each degree ($IPC_{D}$), while the bottom row panels provides a detailed breakdown of individual $IPC_{D}$ values. An enlarged view of the smaller IPC values is shown in the insets in the bottom panels. The values of IPC for up to degree 9 are obtained. The model parameters $(g,\sigma_{s},\sigma_{n})$ are $(0.9,0.3,0.0)$ for (a), $(2.5,2.0,0.0)$ for (b), and $(0.8,0.3,0.1)$ for (c), respectively. While $\sum_{D=1}^{9}IPC_{D}$ almost equals $L$ for (a), it falls significantly short of $L$ for (b) and (c) (a dashed line denotes $\sum_{D=1}^{\infty}IPC_{D}=L$). This is attributed to the violation of the echo state property. Specifically, the RNN for (b) is chaotic, as indicated by a positive maximum conditional Lyapunov exponent [22, 15] ($\lambda_{\rm max}=0.087\pm 0.0067$), while the RNN for (c) is subject to noise. Shaded area represents mean±$\pm$±std for 10 different network and input signal realizations. Simulation time length is $T=10^{5}$.
VII Discussion

### v3

@@ V Neuronal correlations shape the sublinear scaling of memory capacity
Notably, the $n$-th term corresponds exactly to the $n$-th term in the Neumann series expansion, Eq.(IV). Therefore, under the assumption of vanishing neuronal correlations ($C_{\rm nondiag}=O$), the summation terms in Eq.(V) disappear, resulting in $r(L)=1$, and thus, linear scaling. This implies that higher neuronal correlations contribute to the faster decay of $r(L)$ via the summation terms in Eq.(V). In fact, we confirm that $r(L)$ decays more rapidly for hyperparameters that lead to higher neuronal correlation, such as larger input signals ($\tilde{\sigma}_{s}^{2}$), smaller recurrent weights ($g$), and lower noise intensity ($\sigma_{n}^{2}$), as shown in Fig.3.
Importantly, while the slope of the memory capacity curve at the origin, which is proportional to $MC(1)$, reaches its maximum when the reservoir RNN operates near the edge of chaos [14, 15], the extent of the decay of $r(L)$ varies monotonically with the hyperparameters $g$, $\sigma_{s}^{2}$ and $\sigma_{n}^{2}$, irrespective of whether the RNN is near the edge of chaos. This finding suggests a new perspective: instead of proximity to the edge of chaos, neuronal correlation is a novel mechanism that governs the computational ability of reservoir computing.
Figure 3: Analytical solutions for the decay rate of memory capacity as a function of $\alpha$. The activation function is an error function, $\phi(x)=\int_{0}^{x}e^{-\frac{\pi}{4}t^{2}}dt$. Each figure varies a single parameter: (a) input intensity $\tilde{\sigma}_{s}^{2}$, (b) recurrent weight scale $g$, (c) noise intensity $\sigma_{n}^{2}$, while the remaining parameters are fixed ($g=1.2$, $\tilde{\sigma}_{s}^{2}=1.0^{2}$, $\sigma_{n}^{2}=0.5^{2}$). A gradient from darker to lighter gray lines indicates a decreasing level of neuronal correlations. The decay rate for systems with strong neuronal correlations decreases more rapidly.
We have thus far derived the memory capacity under the specific set of scaling assumptions (Eq.(4)). However, practical applications often deviate from these conditions, raising the question of whether neuronal correlations remain a key determinant of sublinear memory capacity growth when these assumptions are relaxed. To explore this, Fig.4(a) examines a broader range of model parameters, including larger input intensities, various noise levels, and different values of $g$. Here, we introduced a half-life of memory capacity growth, $L_{\rm half}$, defined as the value of $L$ at which $r(L)$ equals $0.5$ (Fig.4(a), inset). A lower $L_{\rm half}$ indicates more rapid decay of $r(L)$. The overall strength of neuronal correlations is quantified by the root mean square (RMS) of pairwise Pearson correlation coefficients, $\sqrt{\langle\rho_{ij}^{2}\rangle_{\rm pairwise}}$, where
$\rho_{ij}\equiv\frac{\langle x_{i}x_{j}\rangle-\langle x_{i}\rangle\langle x_{j}\rangle}{\sqrt{\left(\langle x_{i}^{2}\rangle-\langle x_{i}\rangle^{2}\right)\left(\langle x_{j}^{2}\rangle-\langle x_{j}\rangle^{2}\right)}}.$
Note that due to the symmetry of the weight distribution, the time-averaged value $\langle x_{i}\rangle$ almost vanishes for all neurons. We use the RMS value because the symmetry of the weight distribution creates a pairwise distribution of $\rho_{ij}$ values that is symmetric around zero. Consequently, the simple average $\langle\rho_{ij}\rangle_{\rm pairwise}$ is always zero and thus uninformative, whereas the RMS value properly quantifies the typical magnitude of the neuronal correlations.
As shown in Fig.4(a), strong neuronal correlations are indeed closely associated with a more rapid decay of memory capacity growth, with $L_{\rm half}$ being inversely proportional to the level of neuronal correlations. This finding highlights the persistent influence of neuronal correlations even when our original scaling assumptions are relaxed.
Figure 4: Relationship between neuronal correlation and the half-life of memory capacity growth. The x-axis represents the level of neuronal correlations quantified by the root mean square (RMS) of pairwise Pearson correlation coefficients. The y-axis denotes $L_{\rm half}$, the definition of which is graphically shown in the inset of panel (a). (a) Results for RNNs with $\phi=\tanh$. Each data point corresponds to a set of hyperparameters sampled from $g\sim U(0.2,3)$, $\sigma_{s}\sim U(0.1,3)$ and $\sigma_{n}\sim U(0,3)$. Red points indicate parameter regimes beyond our theoretical framework, as determined by $L_{\rm half}$ surpassing the deviation threshold of the theoretical values of MCMC (Fig.2). (b) Results for RNNs with $\phi(x)=\max(0,x)$ (ReLU). Each data point corresponds to a set of hyperparameters sampled from $g\sim U(0.2,\sqrt{2})$, $\sigma_{s}\sim U(0.1,3)$ and $\sigma_{n}\sim U(0,3)$. For both (a) and (b), a total of 100 samples are plotted. Each data point represents the mean value over 10 network and input realizations. Error bars indicate the standard deviation for both $L_{\rm half}$ and $\sqrt{\langle\rho_{ij}^{2}\rangle_{\rm pairwise}}$, but most of the error bars for the latter are too small to be visible behind the data points. Overall, for both activation functions, $L_{\rm half}$ tends to be smaller for stronger neuronal correlations. All simulations use a network size of $N=10000$ and a simulation time of $T=10^{4}$. The transient washout time is $T_{\rm washout}=10^{3}$.
VI Robustness of our findings across other types of RNNs
Thus far, we have examined the scaling behavior of reservoir RNNs with a sigmoid activation function and recurrent weights drawn i.i.d from a Gaussian distribution. We found that the level of neuronal correlations, which is controlled by hyperparameters such as $g$, $\sigma_{s}^{2}$, and $\sigma_{n}^{2}$ in our model, determines the extent to which the growth rate of memory capacity decays. While our model is widely used in both theoretical [14, 15, 27] and practical [16] settings, a natural question arises: can these findings be generalized to other types of RNNs? Here, we hypothesize that our findings on the relationship between neuronal correlations and memory capacity hold more generally. To test this hypothesis, we numerically study three types of RNNs: those with a ReLU activation function, a heavy-tailed distribution of recurrent weights, and reciprocal motifs.
VI.1 RNN with a ReLU activation function
We first study the neuronal correlations and memory capacity of an RNN described by Eq.(1) with $\phi(x)=\max(0,x)$ (ReLU). All parameter settings are the same as those described in Section II, with the exception that $g$ must be lower than $g_{c}=\sqrt{2}$ because the system diverges for $g>g_{c}$. To demonstrate this, we first derive the dynamical mean-field equation for this system in the limit of $N\to\infty$. The derivation is the same as in Appendix A.1 and yields the following equation:
$[x_{i}(t+1)^{2}]=\frac{g^{2}}{2}[x_{i}(t)^{2}]+s(t+1)^{2}+\sigma_{n}^{2}.$ (14)
This self-consistent equation admits a stable solution of $x_{i}(t)^{2}]$ only when $g<g_{c}=\sqrt{2}$. For $g>g_{c}$, the system diverges regardless of the values of $s(t)$ and $\sigma_{n}$.
Our theoretical framework for the analytical derivation of neuronal correlations and memory capacity cannot be applied to an RNN with the ReLU activation function. This limitation arises because our theory utilizes Price’s theorem (Eq.(42)), which cannot be applied to the ReLU function as it is non-smooth at the origin. Therefore, we numerically calculate neuronal correlations and memory capacity, and plot them to examine their relationship.
Following the same procedure as in the previous section, we numerically calculated the strength of neuronal correlations and $L_{\rm half}$. The results, shown in Fig. 4(b), reveal a negative correlation between the level of neuronal correlations and $L_{\rm half}$. Although this correlation is less pronounced than for the sigmoid network (Fig. 4(a)), the finding nonetheless supports our hypothesis that neuronal correlations govern the scaling behavior of memory capacity.
VI.2 RNN with a heavy-tailed distribution of recurrent weights
We now turn to the case of an RNN where the distribution of recurrent weights is heavy-tailed or scale-free. Recent experimental work has revealed that the distribution of synaptic strength in the brain follows a heavy-tailed distribution across a wide range of brain areas [28]. Several theoretical studies have also proposed potential roles for heavy-tailed synapse distributions in brain function [29, 30].
The theoretical analysis of dynamics in large random RNNs, using methods like dynamical mean-field theory or the dynamical cavity method, typically relies on the central limit theorem. This allows the statistical behavior of neurons to be described as a Gaussian process (see Appendix A). However, this assumption breaks down for heavy-tailed weight distributions, where the central limit theorem is not guaranteed to hold. This leads to a fundamental difference in dynamical properties between conventional RNNs and those with heavy-tailed synapses. Although a theoretical framework to derive the dynamical mean-field equations for such RNNs has recently been proposed using the generalized central limit theorem [31], the analysis of neuronal correlations remains challenging. Consequently, we must rely on numerical simulations to investigate the neuronal correlations and memory capacity of these RNNs.
The model examined here is identical to that described in Section II, with the exception that the recurrent weights are i.i.d drawn from a Cauchy distribution [30]:
$p(J_{ij})=\frac{1}{\pi}\frac{\gamma/N}{(\gamma/N)^{2}+J_{ij}^{2}}.$ (15)
Here, the hyperparameter $\gamma>0$ determines the strength of recurrent weights. In the limit of $N\to\infty$, the eigenvalue spectrum of $J$ is unbounded and its spectral radius is infinite for any $\gamma>0$ [31], leading to chaotic dynamics in the RNN regardless of $\gamma$. On the other hand, when $N$ is finite, whether the dynamics are chaotic or stable depends on the specific realization of $J$ [32].
@@ VI.2 RNN with a heavy-tailed distribution of recurrent weights
As demonstrated in Fig.5(a), the level of neuronal correlations is controlled by the parameter $\gamma$, analogous to the role of the parameter $g$ in our canonical Gaussian model. These neuronal correlations are considered to be of $O(1)$ with respect to the network size $N$, as indicated by the nearly constant values of $\sqrt{\langle\rho_{ij}^{2}}\rangle_{\rm pairwise}$ for increasing $N$.
Since we have observed that larger $\gamma$ is associated with weaker neuronal correlations, our hypothesis predicts that the growth rate of memory capacity decays more slowly for larger $\gamma$. This prediction is supported by the results in Fig. 5(b), which shows that the decay rate $r(L)$ decreases more slowly as $\gamma$ increases. Therefore, we conclude that our findings generalize to RNNs with a heavy-tailed distribution of recurrent weights.
VI.3 RNN with reciprocal motifs
The models examined thus far assume that recurrent weights are mutually independent. However, this assumption is not always valid in real-world complex networks. In the brain, for example, certain connectivity motifs are observed more frequently than expected from random connections, which significantly impacts network dynamics [33]. Although neuronal correlations in such networks have yet to be studied, it is reasonable to assume that the prevalence of specific connectivity patterns also affects the strength of neuronal correlations. Therefore, we next investigate whether our findings on the relationship between neuronal correlations and memory capacity generalize to RNNs that incorporate such connectivity motifs.
The RNN model studied here is nearly identical to that described in Sec. II. The primary difference is that the recurrent weight matrix is now defined as $gJ$, where $g>0$ is a gain parameter. The elements of the matrix $J$ are drawn from a Gaussian distribution with zero mean, $1/N$ variance, and a specific correlation structure given by $J_{ij}J_{ji}]_{J}=\eta/N$, following the approach of Marti et al. [34]. The parameter $\eta\in[-1,1]$ controls the degree of (anti-)symmetry of connections, i.e., the prevalence of reciprocal motifs. The model reduces to the canonical random RNN when $\eta=0$, while $\eta=1$ and $\eta=-1$ correspond to fully symmetric and fully anti-symmetric connectivity matrices, respectively. It should be noted that we implement this RNN in a discrete-time model, unlike the original continuous-time formulation by Marti et al. [34], because memory capacity is not yet well-defined for continuous-time systems.
In the absence of inputs and noise, the dynamics of the RNN can be analyzed through the eigenvalue spectrum of the recurrent weight matrix $gJ$. According to the Elliptic Law from random matrix theory, the eigenvalues are uniformly distributed over an elliptical disk in the complex plane, with real and imaginary semi-axes of length, $g(1+\eta)$ and $g(1-\eta)$, respectively [35]. Therefore, the trivial fixed point at zero activity, $\bm{x}=\bm{0}$, is stable if and only if this entire eigenvalue spectrum lies within the unit circle. Otherwise, the system can potentially transition to other dynamical regimes, such as a non-zero fixed point, limit cycle, or chaos. While the phase diagram for the continuous-time version of this model was characterized by Marti et al. [34], the corresponding diagram for our discrete-time model is currently unknown. In fact, our own simulations suggest that its unstable regime contains a rich variety of dynamics, including complex oscillations and chaos (data not shown). Therefore, to ensure a clear interpretation of the results, we confine our numerical simulations to the parameter regime where the zero fixed point is stable. Specifically, we fix $g=0.6$ and vary $\eta$ in the range $2/3<\eta<2/3$, which guarantees $g(1\pm\eta)<1$ is satisfied.
VI.3.1 Autocorrelation function
First, we examine how the autocorrelation function of our model changes as a function of $\eta$. To do this, we numerically calculate the population-averaged autocorrelation function defined as $A(d)=\frac{1}{N}\sum_{i}^{N}\frac{\langle x_{i}(t)x_{i}(t-d)\rangle}{\langle x_{i}(t)^{2}\rangle}$. As shown in Fig. 6(a), an RNN with a larger $|\eta|$ shows a stronger autocorrelation, $|A(d)|$. In contrast, the canonical random model ($\eta=0$) shows a vanishing autocorrelation for all time lags $d\geq 1$. This indicates that a higher density of reciprocal motifs promotes longer timescales in neuronal activity, a finding consistent with continuous-time models [34]. A key difference from the continuous-time model, however, is that $A(d)$ is non-zero only for even time lags and vanishes for odd $d$. Intuitively, this is because a perturbation to neuron $i$ at time $t$ can propagate back to itself at time $t+2$ via a two-step path ($i\to j\to i$). The activity at time $t$ is therefore correlated with the activity at $t+2$ due to the underlying correlation between the synaptic weights $J_{ij}$ and $J_{ji}$. It is possible to derive an analytical expression for $A(d)$ using the dynamical cavity method, but we leave this investigation for future work.
[…]
VI.3.2 Memory function
An RNN with long-timescale activity is expected to be better at retaining inputs from the distant past. To test this prediction, we calculate $M_{d}$ (defined in Eq.(3)) for all $d$, which is sometimes referred to as a memory function. Consistent with our prediction, a larger $|\eta|$ value, which yields a stronger autocorrelation, improves the recall of distant inputs, as shown in Fig. 6(b). However, these same RNNs perform worse at precisely recalling recent inputs, as indicated by their lower $M_{d}$ values for short time lags. This happens because the long intrinsic neuronal timescales impose a slow timescale on the reservoir’s output, preventing it from tracking rapidly changing Gaussian white noise inputs. We believe this trade-off between recent and distant memory is a general principle in reservoir computing. Indeed, a similar effect was reported in a previous study on a plastic spiking neural network reservoir (see Fig.7(a) in [36]). It is important to note that memory capacity, the primary focus of this study, is the sum of $M_{d}$ over all $d$, and thus, a longer neuronal timescale does not necessarily guarantee a higher memory capacity.
VI.3.3 Neuronal correlations
Before assessing memory capacity, we examine whether the parameter $\eta$ affects the strength of neuronal correlations. Fig. 6(c) demonstrates that as $|\eta|$ increases, the strength of neuronal correlations actually decreases. This result may seem counter-intuitive. However, we confirmed that this observation is not a finite-size artifact, as the RMS correlation value remains constant for increasing network sizes $N$. Investigating whether this result is generalizable to other parameter sets and exploring its underlying mechanisms are promising directions for future work, potentially using the dynamical cavity method.
VI.3.4 Scaling behavior of memory capacity
Given our observation that $|\eta|$ determines the level of neuronal correlations, our hypothesis suggests that it would also influence the decay rate of memory capacity. Our numerical simulations corroborate this prediction. As shown in Fig. 6(d), larger values of $|\eta|$, which produce weaker neuronal correlations, lead to a slower decrease in the decay rate $r(L)$. Notably, the value of $\eta$ has no effect on the memory capacity of a single neuron, i.e., $MC(L=1)$, while its impact becomes increasingly significant as the number of readout neurons grows. Therefore, the prevalence of reciprocal motifs affects the computational ability of the network as a whole, rather than that of individual neurons, by modulating the structure of neuronal correlations.
VII Scaling behavior of nonlinear computational capability
@@
While our previous analysis has focused on memory capacity, non-linear computational ability is equally crucial for RC to address complex tasks. As established in previous works, there exists a trade-off between memory capacity and nonlinear computational ability for RC  [13, 27]. Thus, we anticipate that while the growth in memory capacity decelerates as $L$ increases, the system’s capacity for nonlinear computation shows a corresponding acceleration.
@@
Information Processing Capacity (IPC) introduced by Dambre et al. [13] offers a task-independent metric for evaluating the performance of RC, enabling a comprehensive assessment of both linear and non-linear computational capabilities. Since IPC theory is somewhat complicated, we provide detailed explanations in Appendix C.1. Put simply, the IPC for degree $D$, denoted by $IPC_{D}$, represents the reservoir’s ability to approximate $D$-th order polynomial functions of past white noise input signals. The IPC has two key properties. First, by definition, memory capacity is exactly equivalent to $IPC_{1}$, and thus, $IPC_{D}$ for $D\geq 2$ represents non-linear computational ability. Second, it has been proven that the total IPC, defined as $IPC_{\rm total}\equiv\sum_{D=1}^{\infty}IPC_{D}$, equals to the number of readout units, $L$, provided that the reservoir is a fading memory dynamical system [13]. A fading memory dynamical system is defined as a system whose state is uniquely determined solely by its input signals [13]. For example, a reservoir with chaotic dynamics or one subject to noise violates the fading memory property, because its state is determined not only by input signals but also by its initial state or by noise.
@@
Fig.7(a)-(c) illustrate how the values of $IPC_{D}$ of the canonical random RNN, defined in Sec. II, change with increasing $L$ across various hyperparameter combinations ($g$, $\sigma_{s}^{2}$, and $\sigma_{n}^{2}$), as obtained through numerical simulations. Note that, due to the model’s symmetry, $IPC_{D}$ values for all even degrees $D$ are identically zero. As shown, a decline in the growth of the $IPC_{1}$ (equivalent to memory capacity) is accompanied by an increase in the non-linear computational ability (represented by $IPC_{D}$ for $D\geq 3$), reflecting memory-nonlinearity trade-off.
Our novel finding is that, as $L$ increases, the values of higher-degree $IPC_{D}$ are activated sequentially from lower to higher orders, accompanied by their supralinear rises. This distinctive scaling behavior of nonlinear computational ability might be considered a trivial result in RNNs that satisfy the fading memory property, where the total capacity is constrained by $\sum_{D\geq 1}IPC_{D}=L$. In such cases, a sublinear growth in memory capacity ($IPC_{1}$) necessarily implies that the remaining capacity grows supralinearly. However, the significance of our finding lies in its generality. This sequential, supralinear arising of nonlinear computational powers is observed not only in the RNNs that satisfy the fading memory property (Fig. 7(a)), but also in chaotic (Fig. 7(b)) and noisy (Fig. 7(c)) RNNs, where this property breaks down and thus the equality $MC+\sum_{D\geq 2}IPC_{D}=L$ does not hold. Furthermore, the phenomenon holds across various network architectures, including those with a ReLU activation function (Sec. VI.1), heavy-tailed weight distributions (Sec. VI.2), and reciprocal connectivity motifs (Sec. VI.3), as shown in Fig. 7(d)-(f).
@@
As analytical derivation of higher-degree $IPC_{D}$ is beyond our capability, the detailed mechanism underlying these observations remains unclear. However, it is plausible that the neuronal correlation plays a pivotal role, analogous to its influence on memory capacity. This hypothesis is supported by speculation that each $IPC_{D}$ would exhibit linear scaling with $L$ if the effects of neuronal correlations were neglected, which contrasts with the observed supralinear and sequential emergence of $IPC_{D}$ (see Appendix C.2).
[…]
VIII Discussion
In the present study, we investigate how the computational capacity of large random RNN reservoirs scales as increasing the number of readout neurons, $L$. We develop a theoretical framework for deriving memory capacity that can incorporate the effect of neuronal correlations, which has been ignored for analytical simplicity. Our theory reveals that memory capacity scales sublinearly with $L$, and the extent of this sublinearity is significantly influenced by the strength of neuronal correlations, even though those correlations are quite weak. Although our theory is limited to cases where neuronal correlations scale as $O(1/\sqrt{N})$, a constraint imposed by specific assumptions about our model’s hyperparameters (Eq.(4)), our numerical simulations confirm the robustness of this finding across a wider range of hyperparameters and in various network architectures. In addition, we numerically study the scaling behavior of nonlinear computational ability, showing that higher-order computational powers emerge successively, accompanied by their supralinear rises, enabling the reservoir to handle increasingly complex tasks.